OpenStudy (anonymous):

Simplify 4x/9 + x-10/9

OpenStudy (anonymous):

\[\frac{4x}{9}+\frac{x-10}{9}=\frac{4x+x-10}{9}\] is a start

OpenStudy (anonymous):

Awesome, so the final product would be 5x-10/9 right?, since you add 4x and x